Prove or Disprove that the commutative property holds on the matrix multiplication.


1
Expert's answer
2021-10-20T01:17:21-0400

"\displaystyle
\text{Commutative property holds if $MN = NM$, for matrices M and N}\\
\text{Matrix multiplication isn't possible for matrices that aren't square e.g if we have}\\
\text{2 matrices M and N}\\
M_{ij} \cdot N_{jk} \neq N_{jk} \cdot M_{ij}\\
\text{As seen in the right hand side of the expression above the matrix isn't even possible}\\
\text{Hence matrix multiplication is not possible}\\
\text{Also matrix multiplication is not generally possibly for square matrices, to this end}\\
\text{we give a counter-example}\\
\text{Let $ A = \begin{pmatrix}6 & 3\\2 & 5 \end{pmatrix} $ and $B = \begin{pmatrix}-3 & 2\\1 & 5 \end{pmatrix}$}\\
A.B = \begin{pmatrix}-15& 27\\-1 & 29 \end{pmatrix}\\
B.A = \begin{pmatrix}-14 & 1\\16 & 26 \end{pmatrix}\\
\text{Therefore, since $A\cdot B \neq B \cdot A$, hence matrix multiplication isn't commutative}"